# MuSeD: A Multimodal Spanish Dataset for Sexism Detection in Social Media Videos

### v1

@@ Abstract
Sexism is generally defined as prejudice and discrimination based on sex or gender, affecting every sector of society, from social institutions to relationships and individual behavior. Social media platforms amplify the impact of sexism by conveying discriminatory content not only through text but also across multiple modalities, highlighting the critical need for a multimodal approach to the analysis of sexism online. With the rise of social media platforms where users share short videos, sexism is increasingly spreading through video content. Automatically detecting sexism in videos is a challenging task, as it requires analyzing the combination of verbal, audio, and visual elements to identify sexist content. In this study, (1) we introduce MuSeD, a new Multimodal Spanish dataset for Sexism Detection consisting of $\approx$ 11 hours of videos extracted from TikTok and BitChute; (2) we propose an innovative annotation framework for analyzing the contribution of textual and multimodal labels in the classification of sexist and non-sexist content; and (3) we evaluate a range of large language models (LLMs) and multimodal LLMs on the task of sexism detection. We find that visual information plays a key role in labeling sexist content for both humans and models. Models effectively detect explicit sexism; however, they struggle with implicit cases, such as stereotypes—instances where annotators also show low agreement. This highlights the inherent difficulty of the task, as identifying implicit sexism depends on the social and cultural context. [^1] The dataset annotations will be made available for research purposes at https://github.com/lauradegrazia/MuSeD_sexism_detection_videos Warning: This paper contains examples of language and images which may be offensive.
@@ 1 Introduction
The automatic detection of sexism on social platforms is a crucial task because it supports the work of human moderators, who have been shown to develop PTSD and depression (Newton, 2020), and helps prevent the spread and normalization of toxic narratives. It is also a challenging task, particularly when sexism appears in multimodal forms, where multiple modalities contribute to the production of discriminatory content. The identification of sexist content has primarily focused on text, particularly extracting data from X (Rodríguez-Sánchez et al., 2020; Samory et al., 2021; Plaza et al., 2023). Investigating sexism across multiple modalities is an increasingly relevant topic due to the growing diffusion of multimedia content online. The majority of studies in the area of multimodal sexism have focused on the analysis of memes—images that combine pictorial content with overlaid text added on top—and the examination of hate speech in videos (Chhabra & Vishwakarma, 2023). Previous work on identifying sexism in videos has concentrated on analyzing sexism by sex, primarily targeting women (Arcos & Rosso, 2024), while the analysis of sexism based on sex (women), sexual orientation, and gender identity (O’Brien, 2009) remains largely unexplored.
@@
In this work, we aim to address the following research questions: Q1: How can we create a systematic annotation framework to analyze the impact of textual, vocal, and visual modalities in conveying sexist or non-sexist content? Q2: To what extent does visual information play a key role in identifying sexism for both humans and models? Q3: How do model misclassifications relate to human disagreement? Our work makes the following contributions: 1. We introduce MuSeD, a Multimodal Spanish Dataset for Sexism Detection, expanding the definition of sexism to encompass discrimination based on sex, sexual orientation and gender identity. MuSeD includes content that discriminates based on sex, such as portraying women through stereotypes; sexual orientation, by including content that discriminates against non-heterosexual identities; and gender identity, by incorporating content that discriminates against transgender individuals. Table 1 presents representative examples in our dataset. 2. We propose an innovative annotation framework for labeling sexist and non-sexist content, with annotations conducted at different levels: (i) text, (ii) audio, and (iii) video (text, speech, and image), to examine the contribution of visual and multimodal information to identifying sexist and non-sexist content. This is the first work to examine sexism in multimodal data using such a multi-level annotation scheme. 3. We evaluate LLMs and multimodal LLMs on sexism detection using text-only and multimodal inputs, analyzing their alignment with human annotations. We find that visual cues help in labeling sexism, yet models struggle with implicit cases like stereotypes, where annotator agreement is lower. This highlights the challenge of detecting sexism in multimodal content.
Platform Language Target group Annotation of modalities Annotator expertise
@@
Fersini et al. (2019) FB, TW, IG, R eng W Crowd Sourcing
Fersini et al. (2022) TW, R, MGTOW eng W Crowd Sourcing
@@
Plaza et al. (2024) Google image eng, spa W Crowd Sourcing
@@
Arcos & Rosso (2024) TT eng, spa W Trained annotators
MuSeD (Ours) TT, BitChute spa W, LGBTQ+ Moderator, one expert annotator, trained annotators
Table 2: A comparison of existing multimodal datasets for annotating sexism and misogyny online. MuSeD comprises Spanish-language videos from TikTok and BitChute and introduces modality-specific annotations for sexism detection. W: Women, FB: Facebook, TW: Twitter, IG: Instagram, R: Reddit, P: Pinterest, TT: TikTok.
@@ 2.1 Data resources for multimodal sexism
Table 2 compares existing datasets for analyzing sexist and misogynistic content online. We observe that the primary focus of previous work is the analysis of sexism in memes, considering the relationship between text and image. The key differences between our study and previous work are: (i) we broaden the definition of sexism to include discrimination against a target group based on sex, sexual orientation, or gender identity (O’Brien, 2009), whereas previous work primarily focused on sexism based solely on sex; (ii) we select two different platforms as data sources: TikTok, a moderated platform, and BitChute, a low-moderation platform, which enables us to find diverse material; (iii) the annotation task is conducted across different modalities to evaluate the impact of each modality on identifying sexist content. Furthermore, the task involves selecting segments in which annotators identify sexism. This method allows us to precisely identify where sexism occurs, providing a more accurate annotation; (iv) our annotation team includes a moderator expert on sexism in social media and an expert annotator specializing in the annotation of discriminatory content.
@@ 3 MuSeD: A dataset for Multimodal Sexism Detection
MuSeD is a Multimodal Sexism Detection dataset in Spanish. We propose an annotation framework to analyze the impact of each modality (text, speech, and image) on identifying sexist content in videos.
@@ 3.2 Data collection
The data was collected between May and October 2024 from TikTok and BitChute. To gather videos from TikTok, we compiled a list of Spanish hashtags. We did not differentiate between European and Latin American Spanish and included both varieties. The list of Spanish hashtags includes 11 from Plaza et al. (2024), which we significantly expand with 176 additional hashtags to cover discrimination based on sexual orientation and gender identity. The original set is focused primarily on sexism against women, including terms such as #brecha salarial (gender pay gap), #violencia machista (gender-based violence), and #estereotipos de género (gender stereotypes). It also includes general terms such as “mujer” (woman) and “hombre” (man), which, while not explicitly linked to sexism, can be used in sexist contexts. We extended the set to include terms related to sexist content targeting sexual orientation and gender identity such as #lgbt, #binarie (binary), and #ideología de género (gender ideology). The final list consists of 187 hashtags encompassing various topics related to sexuality, gender, feminist discussions and discrimination based on sex and gender. Additionally, we identified a small number of prolific users (Anzovino et al., 2018) who frequently share sexist content—such as relationship advice and recommendations on how to be a “successful” man—as well as other users who share non-sexist content, including discussions on feminist topics and personal stories about experiences of sexist behavior. Based on this search, we collected video URLs associated with the selected hashtags using Apify (Arcos & Rosso, 2024). [^4] https://apify.com We then collected the videos partially using Apify and partially through manual downloads. We ensured that all collected videos had been made publicly available by creators. To obtain the videos from BitChute, we adapted a sample of hashtags from our curated list to use as keywords for the front page search. We then selected a set of videos from the retrieved list and collected them using the BitChutedl software. [^5] https://pypi.org/project/bitchute-dl/ The final dataset includes 400 videos, primarily from TikTok ($\approx$ 90%), while the remaining videos were collected from BitChute.
@@ 3.3 Annotation process
During the annotation process, various biases can arise due to multiple factors such as a lack of demographic diversity among annotators, preconceived stereotypes, and inadequate training (Hovy & Spruit, 2016; Sap et al., 2022). Following the methodology outlined by Zeinert et al. (2021), we employed multiple strategies to mitigate annotator biases. [^8] See the Ethics statement in Section Ethics statement for details. First, we recruited annotators with diverse gender and age backgrounds. [^9] See Table 4 in Appendix B. One of the annotators is an expert in tasks related to the study topic. Second, the annotators underwent training prior to starting the annotation task, which included annotating a sample of 50 items followed by a discussion. Third, we held weekly meetings and, when necessary, updated the annotation guidelines. Each meeting was moderated by an expert on sexism on social media, who led the discussions, particularly on ambiguous cases in which sexism appeared in implicit forms.
@@
In designing our annotation framework, we introduced a novel methodology to classify sexist and non-sexist content across different modalities. The annotation guidelines are included in Appendix A. Our annotation process consisted of three levels: first, annotators labeled the transcript and the OCR texts; [^10] See preprocessing in Section 3.2. second, they annotated the audio; and finally, they annotated the entire video, which included all the modalities. OCR transcriptions were also considered as text because we observed that, in some cases, classification would not be possible without them due to the lack of contextual information. This was particularly relevant in 13 cases with only background music and no identifiable speakers. The group of annotators was split into two teams, each consisting of three people (two who identify as women and one who identifies as a man). Each team annotated half of the dataset in two formats—text and video—while the audio was annotated by the team that did not annotate the text and video. In all three steps, annotators classified the given item (text, audio, or video) as sexist or non-sexist. To do so, they selected the segment they identified as containing sexism.
Content is considered sexist in four main cases: Stereotype: It defines a set of properties that supposely differentiate men and women, based on stereotypical belief (Samory et al., 2021). Inequality: it claims that gender inequalities no longer exist and that the feminist movement is marginalizing the position of men in society (Samory et al., 2021). Discrimination: it discriminates against individuals based on their sexual orientation or gender identity and denigrates the LGBTQ+ community (Chakravarthi et al., 2021), often by accusing them of promoting a “gender ideology”. Objectification: it depicts women as physical objects valued primarily for their utility to others (Szymanski et al., 2011). This type of content is more prevalent in visual form, as images intensify the hypersexualization of female bodies. Table 1 includes examples for each case of sexism. If an item contains a denunciation or a report of a sexist experience, it must be labeled as non-sexist (Chiril et al., 2020).
@@ 3.4 Dataset analysis
Our final dataset, MuSeD, includes 400 videos spanning over $\approx$ 11 hours of content. [^11] See Table 5 in Appendix D for the dataset statistics. The class distribution is balanced with 48.5% of videos labeled as sexist and 51.5% labeled as non-sexist. We observe that the most frequent hashtag is #feminismo (feminism), followed by #hombres (men), #mujeres (women), #machismo (male dominance), and #identidadsexual (sexual identity). [^12] See Figure 3 for the distribution of video lengths and Figure 4 for the frequency of the top 10 hashtags in Appendix D.
@@
The reliability of the annotations was evaluated by analyzing the level of agreement among annotators (Artstein, 2017). To measure IAA, we used Fleiss’ kappa (Fleiss, 1971), observing a consistent increase in agreement when all modalities were included. The agreement increases from 0.719 (Team 1) and 0.717 (Team 2) for text annotation to 0.741 (Team 2) and 0.824 (Team 1) for audio. For video annotation, the IAA reaches 0.834 for Team 1 and 0.853 for Team 2, indicating an improvement from substantial to almost perfect agreement when comparing text and video annotations. In comparison, the study by Arcos & Rosso (2024) reported an IAA of 0.499 for the task of sexism detection, indicating moderate agreement. Based on discussions with the annotators, we identified two main challenges in recognizing sexism: the difficulty of identifying implicit forms of sexism and, in some cases, the contradiction between verbal and non-verbal content. [^13] See Appendix C for examples of the challenging cases.
@@
To evaluate the reliability of annotation agreement, we also selected a sub-sample of the dataset that includes videos extracted from BitChute. We expected the majority of these videos to be marked as sexist, given that BitChute is a platform with low-moderation. We found that, for the annotation task, 87.88% of the videos were marked as sexist. This percentage increased in the video annotation task, where 93.94% of the videos were classified as sexist.
@@ 4.1 Models
4.2 Experimental setup
We evaluate all models in a zero-shot setting by prompting them with a yes/no question about whether the video exhibits sexist content, giving them either the transcript alone or both text and visual information. The system instruction and user prompt are provided in Spanish. Model responses, given in Spanish ( for “Yes” or No for “No”), are then mapped to the corresponding labels (sexist or non-sexist). [^15] Implementation details are included in Appendix E and the prompts in Appendix F.
Evaluation metrics
We evaluate model performance with accuracy, given that the dataset is balanced. [^16] We include F1 scores in Appendix E. Each model is assessed against two types of ground truth labels: the Text Label, which is assigned by annotators based on the text transcript alone, and the Multimodal Label, which is assigned when both text and visual content were available to annotators. With 7% of cases differing between labels, this evaluation enables the analysis of whether text models can predict sexism as perceived in full multimodal context and whether vision-and-language models rely on visual cues that change their predictions with respect to text-based content. Additionally, we include random and majority label baselines.

### v2

@@ Abstract
Sexism is generally defined as prejudice and discrimination based on sex or gender, affecting every sector of society, from social institutions to relationships and individual behavior. Social media platforms amplify the impact of sexism by conveying discriminatory content not only through text but also across multiple modalities, highlighting the critical need for a multimodal approach to the analysis of sexism online. With the rise of social media platforms where users share short videos, sexism is increasingly spreading through video content. Automatically detecting sexism in videos is a challenging task, as it requires analyzing the combination of verbal, audio, and visual elements to identify sexist content. In this study, (1) we introduce MuSeD, a new Multimodal Spanish dataset for Sexism Detection consisting of $\approx$ 11 hours of videos extracted from TikTok and BitChute; (2) we propose an innovative annotation framework for analyzing the contributions of textual, vocal, and visual modalities to the classification of content as either sexist or non-sexist; and (3) we evaluate a range of large language models (LLMs) and multimodal LLMs on the task of sexism detection. We find that visual information plays a key role in labeling sexist content for both humans and models. Models effectively detect explicit sexism; however, they struggle with implicit cases, such as stereotypes—instances where annotators also show low agreement. This highlights the inherent difficulty of the task, as identifying implicit sexism depends on the social and cultural context. [^1] The dataset annotations will be made available for research purposes at https://github.com/lauradegrazia/MuSeD_sexism_detection_videos Warning: This paper contains examples of language and images which may be offensive.
@@ 1 Introduction
The automatic detection of sexism on social platforms is a crucial task because it supports the work of human moderators, who have been shown to develop PTSD and depression (Newton, 2020), and helps prevent the spread and normalization of toxic narratives. It is also a challenging task, particularly when sexism appears in multimodal forms, where multiple modalities contribute to the production of discriminatory content. The identification of sexist content has primarily focused on text, particularly extracting data from X (Rodríguez-Sánchez et al., 2020; Samory et al., 2021; Plaza et al., 2023). Investigating sexism across multiple modalities is an increasingly relevant topic due to the growing proliferation of multimedia content online. The majority of studies in the area of multimodal sexism have focused on analyzing memes—images that combine pictorial content with overlaid text— and examining hate speech in videos (Chhabra & Vishwakarma, 2023). While previous work on identifying sexism in videos has primarily focused on sex-based sexism, particularly targeting women (Arcos & Rosso, 2024), the analysis of sexism that also encompasses sexual orientation and gender identity (O’Brien, 2009) remains largely unexplored.
@@
In this work, we aim to address the following research questions: Q1: How can we create a systematic annotation framework to analyze the impact of textual, vocal, and visual modalities in conveying sexist or non-sexist content? Q2: To what extent does visual information play a key role in identifying sexism for both humans and models? Q3: How do model misclassifications relate to human disagreement? Our work makes the following contributions: 1. We introduce MuSeD, a Multimodal Spanish Dataset for Sexism Detection, expanding the definition of sexism to encompass discrimination based on sex, sexual orientation, and gender identity. MuSeD includes content that discriminates based on sex, such as portraying women through stereotypes; sexual orientation, by including content that discriminates against non-heterosexual identities; and gender identity, by incorporating content that discriminates against transgender individuals. Table 1 presents representative examples in our dataset. 2. We propose an innovative annotation framework for labeling sexist and non-sexist content, with annotations conducted at multiple levels—(i) text, (ii) audio, and (iii) video (text, speech, and image)—to examine the contribution of visual and multimodal information to the identification of content as sexist or non-sexist. This is the first work to examine sexism in multimodal data using such a multi-level annotation scheme. 3. We evaluate LLMs and multimodal LLMs on sexism detection using text-only and multimodal inputs, analyzing their alignment with human annotations. We find that visual cues help in labeling sexism, yet models struggle with implicit cases like stereotypes, where annotator agreement is lower. This highlights the challenge of detecting sexism in multimodal content.
Platform Language
Target
group
Annotation
of modalities
Annotator
expertise
@@
Fersini et al. (2019) FB, TW, IG, R eng W
Crowd Sourcing
Fersini et al. (2022) TW, R, MGTOW eng W
Crowd Sourcing
@@
Plaza et al. (2024) Google image eng, spa W
Crowd Sourcing
@@
Arcos & Rosso (2024) $T$ eng, spa W Trained annotators
MuSeD (Ours) TT, BitChute spa
W,
LGBTQ+
Moderator,
one expert annotator,
trained annotators
Table 2: A comparison of existing multimodal datasets for annotating sexism and misogyny online. MuSeD comprises Spanish-language videos from TikTok and BitChute and introduces modality-specific annotations for sexism detection. W: Women, FB: Facebook, TW: Twitter, IG: Instagram, R: Reddit, P: Pinterest, $T$: TikTok.
@@ 2.1 Data resources for multimodal sexism
Table 2 compares existing datasets for analyzing sexist and misogynistic content online. We observe that the primary focus of previous work is the analysis of sexism in memes, considering the relationship between text and image. The key differences between our study and previous work are: (i) we broaden the definition of sexism to include discrimination against a target group based on sex, sexual orientation, or gender identity (O’Brien, 2009), whereas previous work primarily focused on sexism based solely on sex; (ii) we select two different platforms as data sources: TikTok, a moderated platform, and BitChute, a low-moderation platform, which enables us to find diverse material; (iii) the annotation task is conducted across different modalities to evaluate the impact of each modality on identifying sexist content. Furthermore, the task involves selecting segments in which annotators identify sexism in videos. This method allows us to precisely identify where sexism occurs, providing a more accurate annotation; (iv) our annotation team was moderated by an expert on sexism in social media and included an experienced annotator specializing in the annotation of discriminatory content.
@@ 3 MuSeD: A dataset for Multimodal Sexism Detection
MuSeD is a Multimodal Sexism Detection dataset in Spanish. We propose an annotation framework to analyze the impact of each modality (text, speech, and image) when identifying sexist content in videos.
@@ 3.2 Data collection
The data was collected between May and October 2024 from TikTok and BitChute. To gather videos from TikTok, we compiled a list of Spanish hashtags. We did not differentiate between European and Latin American Spanish and included both varieties. The list of Spanish hashtags includes 11 from Plaza et al. (2024), which we significantly expand with 176 additional hashtags to cover discrimination based on sexual orientation and gender identity. The original set is focused primarily on sexism against women, including terms such as #brecha salarial (gender pay gap), #violencia machista (gender-based violence), and #estereotipos de género (gender stereotypes). It also includes general terms such as “mujer” (woman) and “hombre” (man), which, while not explicitly linked to sexism, can be used in sexist contexts. We extended the set to include terms related to sexist content targeting sexual orientation and gender identity such as #lgbt, #binarie (binary), and #ideología de género (gender ideology). The final list consists of 187 hashtags encompassing various topics related to sexuality, gender, feminist discussions and discrimination based on sex and gender. Additionally, we identified a small number of prolific users (Anzovino et al., 2018) who frequently share sexist content—such as relationship advice and recommendations on how to be a “successful” man—as well as other users who share non-sexist content, including discussions on feminist topics and personal stories about experiences of sexist behavior. Based on this search, we collected video URLs associated with the selected hashtags using Apify (Arcos & Rosso, 2024). [^4] https://apify.com We then collected the videos partially using Apify and partially through manual downloads. We ensured that all collected videos had been made publicly available by the creators. To obtain the videos from BitChute, we adapted a sample of hashtags from our curated list to use as keywords for the front page search. We then selected a set of videos from the retrieved list and collected them using the BitChutedl software. [^5] https://pypi.org/project/bitchute-dl/ The final dataset includes 400 videos, primarily from TikTok ($\approx$ 90%), while the remaining videos were collected from BitChute.
@@ 3.3 Annotation process
During the annotation process, various biases can arise due to multiple factors such as a lack of demographic diversity among annotators, preconceived stereotypes, and inadequate training (Hovy & Spruit, 2016; Sap et al., 2022). Following the methodology outlined by Zeinert et al. (2021), we employed multiple strategies to mitigate annotator biases. [^8] See the Ethics statement in Section Ethics statement for details. First, we recruited annotators with diverse gender and age backgrounds. [^9] See Table 5 in Appendix B. The annotators were sourced through academic channels and all have a background in linguistics. One of the annotators is an expert in tasks related to the study topic. Second, the annotators underwent training prior to starting the annotation task, which included annotating a sample of 50 items followed by a discussion. Third, we held weekly meetings and, when necessary, updated the annotation guidelines. Each meeting was moderated by an expert on sexism on social media, who led the discussions, particularly on ambiguous cases in which sexism appeared in implicit forms. Although all of our annotators were European, we made concerted efforts to alleviate potential bias arising from limited demographic diversity. We ensured that they were familiar with the social and cultural contexts of Spanish-speaking regions, including both European and Latin American contexts. Annotators were expected to fully understand the content of each video, and in cases of dialectical complexity, these instances were collaboratively discussed to reach a shared understanding and maintain consistent annotations.
@@
In designing our annotation framework, we introduced a novel methodology to classify sexist and non-sexist content across different modalities. The annotation guidelines are included in Appendix A. Our annotation process consisted of three levels: first, annotators labeled the transcript and the OCR texts; [^10] See preprocessing in Section 3.2. second, they annotated the audio; and finally, they annotated the entire video, which included all the modalities. OCR transcriptions were also considered as text because we observed that, in some cases, classification would not be possible without them due to the lack of contextual information. This was particularly relevant in 13 cases with only background music and no identifiable speakers. The group of annotators was split into two teams, each consisting of three people (two who identify as women and one who identifies as a man). Each team annotated half of the dataset in two formats—text and video—while the audio was annotated by the team that did not annotate the text and video. In all three steps, annotators classified the given item (text, audio, or video) as sexist or non-sexist, using Label Studio as a platform. [^11] https://labelstud.io. Moreover, for the video annotation task, they selected the segment they identified as containing sexism.
Content is considered sexist in four main cases. Stereotype: it defines a set of properties that supposely differentiate men and women, based on stereotypical belief (Samory et al., 2021); Inequality: it claims that gender inequalities no longer exist and that the feminist movement is marginalizing the position of men in society (Samory et al., 2021); Discrimination: it discriminates against individuals based on their sexual orientation or gender identity and denigrates the LGBTQ+ community (Chakravarthi et al., 2021), often by accusing them of promoting a “gender ideology”; Objectification: it depicts women as physical objects valued primarily for their utility to others (Szymanski et al., 2011). This type of content is more prevalent in visual form, as images intensify the hypersexualization of female bodies. Table 1 includes examples for each case of sexism. If an item contains a denunciation or a report of a sexist experience, it must be labeled as non-sexist (Chiril et al., 2020).
@@ 3.4 Dataset analysis
Our final dataset, MuSeD, includes 400 videos spanning over $\approx$ 11 hours of content. [^12] See Table 6 in Appendix D for the dataset statistics. We observe that the most frequent hashtag is #feminismo (feminism), followed by #hombres (men), #mujeres (women), #machismo (male dominance), and #identidadsexual (sexual identity). [^13] See Figure 4 for the distribution of video lengths and Figure 4 for the frequency of the top 10 hashtags in Appendix D. The class distribution is balanced with 48.5% of videos labeled as sexist and 51.5% labeled as non-sexist. [^14] See Table 7 for the distribution of sexist and non-sexist videos on TikTok and BitChute. We conduct a deeper analysis of the sexist videos by examining how many were categorized into the four cases of sexism described in Section 3.3. The majority of these videos were labeled as Stereotype (56.2%), followed by Inequality (39.2%), Discrimination (15.5%), and Objectification (3.6%).
@@
The reliability of the annotations was evaluated by analyzing the level of agreement among annotators (Artstein, 2017). To measure IAA, we used Fleiss’ kappa (Fleiss, 1971), observing a consistent increase in agreement when all modalities were included. The agreement increases from 0.719 (Team 1) and 0.717 (Team 2) for text annotation to 0.741 (Team 2) and 0.824 (Team 1) for audio. For video annotation, the IAA reaches 0.834 for Team 1 and 0.853 for Team 2, indicating an improvement from substantial to almost perfect agreement (Landis & Koch, 1977; Artstein & Poesio, 2008) when comparing text and video annotations. In comparison, the study by Arcos & Rosso (2024) reported an IAA of 0.499 for the task of sexism detection, indicating moderate agreement. Based on discussions with the annotators, we identified two main challenges in recognizing sexism: the difficulty of identifying implicit forms of sexism and, in some cases, the contradiction between verbal and non-verbal content. [^15] See Appendix C for examples of the challenging cases.
@@
To evaluate the reliability of annotation agreement, we also selected a sub-sample of the dataset that includes videos extracted from BitChute. We expected the majority of these videos to be marked as sexist, given that BitChute is a platform with low-moderation. We found that, in the text annotation task, 87.88% of the videos were marked as sexist. This percentage increased in the video annotation task, where 93.94% of the videos were classified as sexist.
@@ 4.1 Models
4.2 Experimental setup
We evaluate all models in a zero-shot setting by prompting them with a yes/no question about whether the video exhibits sexist content, giving them either the transcript alone or both text and visual information. The system instruction and user prompt are provided in Spanish. Model responses, given in Spanish ( for “Yes” or No for “No”), are then mapped to the corresponding labels (sexist or non-sexist). [^17] Implementation details are included in Appendix E and the prompts in Appendix F. We also conducted additional experiments using an extended, definition-based prompt [^18] The extended prompts are included in Appendix F.2. that explicitly includes the definition of sexist content, as detailed in our guidelines in Appendix A.
Evaluation metrics
We evaluate model performance with accuracy, given that the dataset is balanced. [^19] We include F1 scores in Appendix E. Each model is assessed against two types of ground truth labels: the Text Label, which is assigned by annotators based on the text transcript alone, and the Multimodal Label, which is assigned when both text and visual content were available to annotators. With 7% of cases differing between labels, this evaluation enables the analysis of whether text models can predict sexism as perceived in a full multimodal context and whether vision-and-language models rely on visual cues that change their predictions with respect to text-based content. Additionally, we include random and majority label baselines.